OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (3, 4) and (5, 16)?

Mehek (mehek14):

to find the slope \(\dfrac{16-4}{5-3}=\dfrac{12}{2}=6\)

Mehek (mehek14):

so your slope is 6

Mehek (mehek14):

right now your equation is \(y=6x+b\)

Mehek (mehek14):

b is the y-intercept which we are going to find

OpenStudy (anonymous):

Solve it, and I'll tell you if you're right

Mehek (mehek14):

put the ordered pair 3,4 into the equation \(4=6*3+b\\6*3=18\\4-18=-14\\-14=b\)

Mehek (mehek14):

so the slope is 6 and the y-int is -14 put that all together and you get \(y=6x-14\)

OpenStudy (anonymous):

@Mehek14 is correct